Calculate the wavelength, in metres, of a photon capable of exciting an electron in He+ from the ground state to n=4... Someone please tell me the right answer urgently :( ?
1
Expert's answer
2012-11-06T07:35:54-0500
Energy levels for the hydrogen-like He+ are given by
E(n) = -Z2*R/n2, where Z = 2;

Energy of an absorbed photon between the levels n=1 and m=4
E(m,n) = Z2*R*(1/n2 - 1/m2)

We will take R = 10973731.6 m-1, Rydberg constant for H, in reciprocal meters.

Photon wavelength
lambda = h*c/E = 1/Z2*R*(1/n2 - 1/m2)) = 1/(22*10,973,731.6 m-1*(1/12 - 1/42)) = 3.746*10-8 m

Answer: 3.746*10-8 meters